### v1

# Strategic resource allocation in memory encoding: An efficiency principle shaping language processing

Abstract
How is the limited capacity of working memory efficiently used to support human linguistic behaviors? In this paper, we investigate strategic resource allocation as an efficiency principle for memory encoding in sentence processing. The idea is that working memory resources are dynamically and strategically allocated to prioritize novel and unexpected information, enhancing their representations to make them less susceptible to memory decay and interference. Theoretically, from a resource-rational perspective, we argue that this efficiency principle naturally arises from two functional assumptions about working memory, namely, its limited capacity and its noisy representation. Empirically, through naturalistic corpus data, we find converging evidence for strategic resource allocation in the context of dependency locality from both the production and the comprehension side, where non-local dependencies with less predictable antecedents are associated with reduced locality effect. However, our results also reveal considerable cross-linguistic variability, highlighting the need for a closer examination of how strategic resource allocation, as a universal efficiency principle, interacts with language-specific phrase structures.
Keywords: strategic resource allocation; working memory efficiency; dependency locality; cross-linguistic; sentence production; sentence comprehension; naturalistic data
1 Introduction
Language processing in humans relies on working memory, a cognitive module known for its limited capacity to retain information (Baddeley, 1992; Just & Carpenter, 1992; Fedorenko et al., 2013). A perceived linguistic signal, under this limitation, is at the risk of being lost, as it is rapidly overwhelmed by the continual torrent of new input (Christiansen and Chater, 2016). Meanwhile, language use seems effortless, with sophisticated linguistic representations being dynamically encoded and decoded within milliseconds. This dual nature of working memory raises the question: how is the limited capacity of working memory efficiently used to support human linguistic behaviors?
In this paper, we argue for strategic resource allocation as an efficiency principle for memory encoding in sentence processing: working memory resources are dynamically and strategically allocated to prioritize novel and unexpected information. We argue that this efficiency principle, as a resource-rational theory (Lieder & Griffiths, 2020; Lewis et al., 2014; Gershman et al., 2015), naturally arises from two functional assumptions about working memory: its capacity is limited, and its representations are noisy. To examine this efficiency principle, we report three studies using naturalistic corpus data, where we demonstrate an interaction between strategic resource allocation and the locality effect in processing non-local syntactic dependencies.
2 Background
We first present the theoretical justification and empirical support for our proposal of strategic resource allocation, drawing from the literature on sentence processing and visual working memory. We then review some critical findings about the dependency locality effect, an empirical domain through the lens of which we examined our proposal.
2.1 Strategic Resource Allocation
2.1.1 Theoretical Justification
We argue that working memory resources are strategically allocated in a way that prioritizes novel and unexpected information given the context, an efficiency principle that we refer to as strategic resource allocation. This principle is in line with the resource-rational analysis of human mind (Lieder & Griffiths, 2020; Lewis et al., 2014; Gershman et al., 2015). Grounded in the bounded-rational approach to cognition (Anderson, 1990; Simon, 1955), resource-rational analysis aims to integrate the functional goals of a computational problem into the structural constraints of the underpinning cognitive architectures, providing a linkage between the computational-level and the algorithmic-level theories (Marr, 1982). In other words, instead of looking for an unbounded optimization, resource-rational analysis seeks to explain human behaviors under bounded rationality, that is, to identify an optimal solution that strikes the balance between maximizing the functional utility and adhering to the structural constraints of the cognitive system.
The resource-rational explanation for strategic resource allocation is rooted in two widely held functional assumptions about working memory. First, the capacity of working memory is assumed to be limited. Although the exact nature of this limitation is still under debate, recent models in some non-linguistic domains have shifted from a discrete slot representation (Miller, 1956; Cowan, 2001; Luck & Vogel, 1997; Pashler, 1988) towards a continuous resource-based representation, where the limited capacity can be flexibly allocated across the encoded information (Ma et al., 2014; van den Berg et al., 2012; Brady et al., 2016; Brady and Tenenbaum, 2013; van den Berg & Ma, 2018; Sims et al., 2012; Sims, 2016; Jakob & Gershman, 2023; Bates and Jacobs, 2020).
Figure 1: Conceptual framework of strategic resource allocation in memory encoding. Different linguistic contexts yield different prior distributions for the input $w^{*}$, one being predictable and the other unpredictable. When encoding $w^{*}$, allocating more memory resources results in sharper encoding with less uncertainty, such that more probability mass is concentrated around the true input. Upon retrieval, although high surprisal $w^{*}$ is less likely to be successfully reconstructed, the one encoded with higher resources is reconstructed with less error than the one encoded with lower resources.
Second, memory representation is full of noise and uncertainties, with unpredictable corruption of the veridical forms of sensory input, resulting in distorted representations that undermine behavioral performance such as inaccurate recall and comprehension (Ma et al., 2014; Gibson et al., 2013; Levy, 2008b; Brady et al., 2024). Given the noisy input, the true state of a past input is assumed to be inaccessible, and memory recall is often implemented as an inferential process from uncertainties using the statistical structure of long-term knowledge, a process often mathematically characterized by Bayes’ rule (Futrell, Gibson, and Levy, 2020; Gibson et al., 2013; Levy, 2008b; Ryskin et al., 2021; Bays et al., 2024):
$p(w\mid r)\propto p_{M}(r\mid w)p(w).$ (1)
Specifically, as shown in Figure 1, a past input unit $w^{*}$ is first encoded as a distribution over representations $r$ over all words with uncertainty. Later in retrieval, the input unit is reconstructed from this noisy memory representation $r$ using prior knowledge $p(w)$. In this process, units that are more probable in the prior are more likely to be accurately reconstructed, resulting in higher posterior recall accuracy.
The two assumptions above then naturally give rise to the following optimization problem: how to maximize memory accuracy under the constraint of limited resources? We argue that strategic resource allocation can be seen as a rational solution to this optimization problem given the reconstruction mechanism described above. Importantly, we assume that the amount of memory resources allocated to a unit is proportional to the precision of the encoded distribution. That is, as illustrated in Figure 1, more resources allocated to encode $w^{*}$ results in a sharper distribution with less uncertainty, such that more probability mass is concentrated around the true input $w^{*}$.
Now, consider a hypothetical situation in which the memory resources are uniformly distributed across all linguistic units regardless of their surprisal. Under this uniform distribution, each unit will receive an identical encoding distribution. However, due to the differences in reconstructability with regard to prior surprisal, the same encoding distribution would yield different retrieval accuracy, disproportionally exerting impact on high surprisal units, reducing their posterior accuracy more significantly than low surprisal units. Therefore, a uniform distribution of memory resources is not the most efficient way to go, and leaves substantial room for improving the overall posterior accuracy.
An alternative approach, which is the one we explore in the current study, is to allocate resources strategically based on a unit’s predictability from context. As mentioned above, we propose that high surprisal linguistic units a priori are prioritized in working memory, receiving more resources than low surprisal ones. That means the prioritized, more surprising units will receive sharper encoding, resulting in less erroneous retrieval. This asymmetric allocation of memory resources is more efficient than the uniform strategy described earlier, in the sense that it achieves higher accuracy across inputs. By sacrificing a slight reduction in posterior retrieval accuracy for low surprisal units, significant gains can be achieved for high surprisal ones by preventing these irrecontructable units from being distorted in the first place. Put simply, when only a limited number of linguistic units can be encoded with minimal distortion, it is more important to encode the more surprising and less reconstructable ones.
Beyond memory, strategic resource allocation aligns with theories such as predictive coding, free energy principle, and implicit learning. At the neural level, the predictive coding mechanism (Murray et al., 2002; Rao & Ballard, 1999; Aitchison and Lengyel, 2017; Blank and Davis, 2016; Sohoglu & Davis, 2020, 2016; Gagnepain et al., 2012) and the free energy principle (Friston, 2005, 2010; Gershman, 2019) hold that the brain seeks to minimize its prediction error, or surprise, as a way to optimize its internal model of the external environment. This principle is implemented by encoding prediction errors rather than the raw sensory input in neural signals. At the behavioral level, implicit learning theories often hold that learning is error-driven, with considerable empirical support that larger prediction errors lead to greater learning effect (Chang et al., 2006; Scheepers, 2003; Hartsuiker & Kolk, 1998; Bock, 1986; Ferreira, 2003; Jaeger & Snider, 2013; Xu & Futrell, 2024; Rumelhart et al., 1986; Elman, 1990; Courville et al., 2006; Wagner & Rescorla, 1972). Taken together, all these theories delivered a similar implication for our proposal in the domain of working memory. That is, when predictions conflict with the actual perceptual input (that is, when there is high surprisal input), it signals the need for comprehenders to update their mental model in order to make more accurate predictions in the future. Given this critical role of more surprising linguistic units in refining the mental model, more cognitive resources should be allocated to them.
2.1.2 Empirical Support
A dynamic similar to strategic resource allocation is observed in the resource-rational model of sentence processing by Hahn et al. (2022). Grounded in the framework of lossy-context surprisal (Futrell, Gibson, and Levy, 2020), their model involves a contextual representation that represents only those words that are most useful for a downstream next-word prediction task. Their model predicts that function words, which are mostly predictable from the linguistic context, are more likely to undergo decay. In fact, our proposal of strategic resource allocation and lossy-context surprisal theory form two sides of the same coin in many aspects. We will discussion the relationship between these two theories in more details in General Discussion.
Studies focusing on the retrieval mechanism in working memory find that linguistic units of higher semantic complexity can be more easily retrieved in later stages of processing despite the initial encoding difficulty, implicating an enhanced accessibility for informative content from the model-theoretic perspective of informativity (Hofmeister, 2011; Hofmeister & Vasishth, 2014; Karimi & Ferreira, 2016; Troyer et al., 2016). However, the exact cognitive underpinning for this empirical observation is still debatable (Hofmeister & Vasishth, 2014; Karimi et al., 2023), and there is lack of clear empirical evidence for whether this effect can be extended to linguistic units of lower predictability (Hofmeister, 2011), a measure that is more consistent with the information-theoretic view of informativity (Shannon, 1948). In spite of these unsettled issues, as a preliminary evidence from the existing literature, the effect of facilitated retrieval for semantically complex units aligns with our rational account outlined above, in the sense that the enhanced accessibility associated with informative units results from the prioritized resources allocated to their encoding.
More recently, strategic resource allocation is more directly examined by Xu & Futrell (2025) through the lens of the agreement attraction effect in English. As shown below, even though the sentences in (1) are ungrammatical in English due to the mismatch of number feature between the subject head noun and the main verb, they are often perceived grammatical by native speakers due to the interference from the distractor noun in between, which shares the number feature with the ungrammatical main verb. In Xu & Futrell (2025), by manipulating the surprisal of the subject head noun through a prenominal adjective, they find that, compared to more surprising subject head nouns (e.g., cute monster), less surprising ones (e.g., evil monster) lead to stronger agreement attraction effect, such that the processing of the main verb is less susceptible to the interference from the distractor noun. They interpret the result as evidence for an enhanced memory representation of more surprising linguistic units against memory interference.
  1. (1)
    1. a.
      *The evil monster who chased the kids seemingly were gone before the sunset. [low surprisal]
    2. b.
      *The cute monster who chased the kids seemingly were gone before the sunset. [high surprisal]
In visual working memory, statistical regularities in long-term knowledge have been shown to shape memory performance. Despite the fact that items more consistent with prior knowledge are easier to be encoded, with lower neural activity and enhanced behavioral performance (Bates and Jacobs, 2020; Blalock, 2015; Jackson & Raymond, 2008; Xie & Zhang, 2017; Girshick et al., 2011), some recent studies indeed observe that, in later stages of processing, these familiar items are de-prioritized to save more resources for the processing of novel ones (Kowialiewski et al., 2022; Bruning and Lewis-Peacock, 2020; Hedayati et al., 2022; Brady et al., 2024). For example, in a delayed-estimation task, Bruning and Lewis-Peacock (2020) ask participants to first memorize and then recall the exact locations of six colored balls on a circle after a brief delay. Before the task, a sub-area on the circle has been previously illustrated to certainly contain the ball with a specific color (e.g., red ball) as a prior information. Their critical finding is that colors not included in the prior information (e.g., non-red balls) have lower recall accuracy when positioned closer to that sub-area, suggesting that memory resources have been shifted away from the prior area to prioritize other areas where novel information is more likely to appear.
2.2 Dependency Locality
The empirical focus of this paper to exmaine the principle of strategic resource allocation is the locality effect in sentence processing, which has been considered a representative example of the efficient use of working memory resources. Consider the sentence pair in (2). In (2a), codependents in the subject–verb dependency are adjacent to each other, whereas in (2b), there is additional linguistic material in between:
  1. (2)
    1. a.
      The monster approached the princess…
    2. b.
      The monster who stayed in the tower approached the princess…
The Dependency Locality Theory (DLT) (Gibson, 1998, 2000) holds that the formation of the non-local structures is constrained by the limited capacity of working memory. Specifically, as dependency distance increases, there is a higher memory cost to store the incomplete dependency as well as a higher integration cost to compute the new structural representation when the later co-dependent is encountered. In support of the DLT, increased processing difficulty is often associated with structures that have longer dependency distance (e.g., Miller & Isard, 1964; Yngve, 1960; Grodner & Gibson, 2005; Bartek et al., 2011; Ford, 1983; King & Just, 1991; Gordon et al., 2001; Traxler et al., 2002). [^1] There is actually an anti-locality effect often found in some head-final dependencies, which is considered to be better explained by an expectation-based mechanism (Vasishth & Lewis, 2006; Nakatani & Gibson, 2010; Konieczny, 2000; Levy & Keller, 2013). Similarly, in the resolution of structural ambiguity where a constituent has multiple potential attachment sites, there is a tendency for comprehenders to prefer the structure with local attachment over others (Frazier and Fodor, 1978; Pearlmutter & Gibson, 2001; Gibson et al., 1996).
Due to the memory constraint involved in processing non-local dependencies, an efficiency principle for language structure should be that linguistic units connected in a syntactic dependency tend to stay close in linear order. This locality principle is evidenced by cross-linguistic word-order patterns (Hawkins, 1990, 1994, 2004; Ferrer-i-Cancho, 2004; H. Liu, 2008; Futrell et al., 2015; Futrell, Levy, and Gibson, 2020; Temperley & Gildea, 2018; Z. Liu, 2021; Z. Liu & Wulff, 2023) (cf. Z. Liu, 2020), and has been argued to explain typological patterns such as the consistency in head direction, the contiguity of constituents, and the asymmetry of short-before-long versus long-before-short between head-initial and head-final languages (Futrell, Levy, and Gibson, 2020; Hawkins, 2004, 1994).
More recently, some studies propose a generalization from dependency locality to information locality, where any pair of linguistic units with high co-occurrence statistics, no matter whether they are in the same syntactic dependency or not, should stay close in linear order (Futrell, 2019; Futrell, Gibson, and Levy, 2020; Hahn et al., 2021; Hahn & Xu, 2022). Compared to previous work, these studies highlight the role of predictive processing, pointing out an interaction between the memory-based and the expectation-based mechanisms. Specifically, under the framework of Surprisal Theory (Levy, 2008a; Hale, 2001), the processing difficulty of a linguistic unit is proportional to how well it is predictable from the memory representation of the past input, which is prone to memory loss and distortion. The locality effect, as an efficient use of working memory, suggests that linguistic units carrying the most relevant information to predict the current one should stay in the recent past before they are forgotten.
These locality principles depend on the precise nature of working memory. Therefore, beyond the general capacity-based constraint proposed by the DLT, it remains an open question how far this efficiency account can go with more and more realistic and detailed characterization of the nature of working memory constraints. Moreover, the existing discussion in the literature rarely addresses efficiency in processing per se. In other words, it is possible that memory limitations make language users not only actively choose a sentence form that is easier to process, but also develop an efficient processing strategy to better handle the information they passively receive.
2.3 The Current Study
Figure 2: Empirical prediction of strategic resource allocation in dependency locality
We examine strategic resource allocation in the context of dependency locality through naturalistic corpus data. If working memory resources are indeed dynamically and strategically allocated such that novel and unexpected information is prioritized, we predict that antecedents (i.e., left codependents) that are more surprising should receive sharper encoding with less uncertainty. The consequence of this is that memory for more surprising antecedents is enhanced, making their representations less susceptible to memory decay and interference before they need to be re-accessed at the other side of the dependency. Therefore, as illustrated in Figure 2, more surprising antecedents should be able to tolerate longer dependency length, resulting in a reduced locality effect. We approach this prediction from both production (Study 1) and comprehension (Study 2a and 2b).
There are two terminological clarifications. First, we adopt a relatively broad interpretation of the term memory encoding in this article, focusing on the representational aspect of working memory mechanisms. Second, the term resources refers to any quantity that is limited and costly to use for better cognitive performance. Given the ongoing debate about the exact nature of working memory resources (Ma et al., 2014; Bays et al., 2024), we choose to restrict the use of this term to its abstract sense.
To preview our results, we find converging evidence from both production and comprehension that unexpected information is encoded with enhanced robustness against decay and interference. In Study 1, which focuses on production data, we observe that more surprising antecedents are associated with longer dependency lengths, an effect that is not reducible to a simple frequency effect. Moreover, the effect mostly exists within Indo-European and head-initial languages in our analysis, and is more consistent for subject relations. In Study 2a and 2b, examining comprehension data from English reading-time corpora, we find a reduced locality effect at the retrieval site for more surprising antecedents. Consistent with Study 1, this effect is more pronounced in subject relations and is observed more reliably in the self-paced reading corpus (Study 2a) than in the eye-tracking corpus (Study 2b).
3 Study 1: Production Side
We first examine strategic resource allocation in dependency locality in production. We predict that in production, the pressure to minimize dependency length can be relaxed when the antecedent contains novel and unexpected information. Consider the subject–verb dependency in the sentences below in (3):
  1. (3)
    1. a.
      The evil monster in the tower approached
    2. b.
      The cute monster who stayed in the tower near the castle approached
The subject “the cute monster” in (3b) is more surprising compared to “the evil monster” in (3a). According to our hypothesis, the unpredictable “cute monster” should be prioritized with more memory resources for encoding, and therefore is more capable of resisting the interference or decay introduced by the intervening material before the verb. As a result, compared to (3a), the less predictable antecedent in (3b) is able to tolerate more intervening material before being re-accessed at the retrieval site (i.e., the right codependent), leading to longer dependency length. We measure the predictability of word $w$ at position $t$ as surprisal $S_{t}$:
$S_{t}\equiv-\log p\left(w_{t}\mid w_{<t}\right),$ (2)
which is the negative log likelihood of the word $w_{t}$ given its preceding context $w_{<t}$. The higher the surprisal, the less predictable a word is. Therefore, we predict a positive correlation between antecedent surprisal and dependency length $L$ in production data.
Besides antecedent surprisal, we also examined the role of antecedent frequency in shaping memory allocation. On the one hand, these two quantities are highly correlated, in that low frequency words are also unpredictable in general, thus yielding higher surprisal. However, on the other hand, compared to surprisal, frequency as a unigram probability does not contain any information from the context. By comparing the effects of antecedent surprisal and antecedent frequency, we aim to look into to what extent the contextual information contributes to the efficiency strategy of working memory encoding. We expect less frequent antecedents to associate with longer dependency length.
3.1 Method
3.1.1 Data
We used the corpora of 10 languages taken from Universal Dependencies (UD) release 2.11 (Nivre et al., 2020), as summarized in Table 1, with the aim to cover a wide variety of typological configurations (e.g., head-initial vs. head-final; free vs. rigid word order). [^2] The original Russian corpus has over 1.2M tokens with over 600 documents; we randomly sampled 300 documents from the original corpus in our analysis in order to save on computational power. An illustration of UD annotations is shown in (4), where each arc represents a dependency whose direction is from the head to the dependent. [^3] In our analysis, antecedent is defined as the left codependent of a dependency, and the retrieval site is always considered the right codependent, although as seen in (4) the direction of a dependency can either go from the left codependent to the right or the other way around.
  1. (4)
    Example of UD annotation:
    {dependency}
    [arc edge, arc angle=85, text only label, label style=above] {deptext}[column sep=1.2em] Isaac & Newton & left & a & note & to & Einstein.
[…]
Compared to the Surface Syntactic Universal Dependencies (SUD) (Gerdes et al., 2018), which is another major project of dependency corpora, the UD annotation scheme is content-word-oriented. That is, UD always labels content words as the head of a unit. As a result, UD favors lexical heads rather than functional heads in cases like adpositions, subordinating conjunctions, auxiliaries, and copulas. For the sentence above in (4), for example, SUD annotates the oblique relation as from the head “left” to the preposition “to” rather than to “Einstein.” In the current work, we chose to use UD corpora since memory processes are more sensitive to content words rather than function words (Gibson, 1998; Grodner & Gibson, 2005).
Some UD corpora consist of out-of-context independent sentences, while others are organized document by document, which provide longer and enriched discourse context for each token. This difference may influence our surprisal estimates, which are sensitive to the preceding context of each token. We extracted all the dependencies of a sentences annotated in the UD corpora. All the UD corpora we used, except for Amharic, have a pre-defined split into training, dev, and test sets. We only used the pre-defined training sets since they already have decent sample size.
{adjustwidth}
-0.2in-0.2in Language Corpus Genre # All # Subj # Obj Danish DDT (Johannsen et al., 2015) sent 45,976 4,203 3,963 English GUM (Zeldes, 2017) doc 89,947 7,881 7,296 German GSD (McDonald et al., 2013) sent 155,480 9,602 8,474 Italian ISDT (Bosco et al., 2013) doc 208,939 10,323 11,735 Japanese GSD (Tanaka et al., 2016) sent 113,771 5,005 4,018 Korean Kaist (Chun et al., 2018) doc 154,609 9,855 24,690 Mandarin GSDSimp (Nivre et al., 2020) sent 63,456 5,538 7,576 Russian SynTagRus (Droganova et al., 2018) doc 329,745 32,822 25,065 Spanish AnCora (Taulé et al., 2008) doc 333,728 21,472 31,143 Turkish BOUN (Marşan et al., 2022) sent 45,914 3,861 4,680
Table 1: Dependency corpora used in Study 1. ‘Genre’ refers to whether the texts in the corpus are organized as independent sentences (‘sent’), or as documents with larger coherent discourse size (‘doc’). ‘# All’ indicates the number of all the dependencies after data exclusion. ‘# Subj’ is a subset of ‘# All’ and indicates the number of dependencies with subject relations. ‘# Obj’ indicates the number of dependencies with object relations.
3.1.2 Estimating Token Surprisal
In this work, to ensure that the results are not the artifact of a specific language model, we generated surprisal measures from both the GPT-3 base (text-davinci-001; Brown et al., 2020) and the mGPT language models (Shliazhko et al., 2024), both being trained on multilingual data. For each token $w$ in the dependency corpora, we obtained its surprisal $-\log p(w_{t}\mid w_{<t})$ given the preceding context from both models. We used the maximally allowed context window in the corresponding document or sentence. It is worth noting that Mandarin Chinese, unfortunately, is not supported by mGPT. Therefore, we only report the results with GPT-3 surprisal for Mandarin.
Contemporary large language models (LLMs) implemented with artificial neural networks provide state-of-the-art probabilistic measures of linguistic sequences and next-word predictions for the approximation of human predictive processing in psycholinguistics research (Wilcox et al., 2020, 2023; Xu et al., 2023; Shain et al., 2024). Empirically, the surprisal generated from LLMs highly correlates with human language processing difficulty indexed by both behavioral and neural responses (Goodkind & Bicknell, 2018; Hao et al., 2020; Shain et al., 2024; Hoover et al., 2023; Schrimpf et al., 2021; Li & Ettinger, 2023; Wilcox et al., 2023; Xu et al., 2023; Hu et al., 2020).
3.1.3 Measuring Dependency Length
We did the analysis with two different measures of dependency length $L$. The first measure is an orthographic one $L_{\mathrm{O}}$, which is the number of words between the codependents of a dependency. The second measure is an information-theoretic one $L_{\mathrm{I}}$, which sums up the surprisal of all the words from $w_{i}$ to $w_{i+N}$:
$\begin{split}L_{\mathrm{I}}&=-\log p(w_{i\dots i+N}\mid w_{<i})\\ &=-\sum^{i+N}_{j=i}\log p(w_{j}\mid w_{<j}).\end{split}$ (3)
We used these two measures because different words presumably induce memory interference to different extents. For example, compared to a content word that marks a discourse referent, a function word such as a determiner is way less informative, and may require much smaller memory load, thus inducing weaker memory interference (Gibson, 1998; Grodner & Gibson, 2005). Compared to the orthographic $L_{\mathrm{O}}$, which treats all the words in the same way, the information-theoretic $L_{\mathrm{I}}$ may better capture the above-mentioned variability across different words (Hahn et al., 2021). [^4] A potential problem with the information-theoretic $L_{\mathrm{I}}$ lies in the dual role attributed to surprisal: it has been theorized as being proportional both to the allocated memory resources and to memory cost. Although intuitively more memory resources allocated may induce higher memory cost as well, we acknowledge that the extent to which these two concepts can be treated as interchangeable remains a debatable question.
3.1.4 Data Transformation and Exclusion
Constructions such as foreign phrases, multi-word proper names, and fixed expressions are annotated as flat structures in UD corpora. We merged flat structures such that the surprisal of the whole structure is the sum of all its components, and that the first word in the flat structure is treated as the head when calculating the length of a dependency involving a flat structure. For example, the subject–verb dependency in (4) involves a flat structure in the subject position. The antecedent surprisal for this dependency is thus the sum of surprisal over both words “Isaac Newton,” and the dependency length by word counts is one since the first word “Isaac” is one word away from the verb. We excluded sentences that are less than five-word long, since sentences that are too short may have limited room for the dependency length to vary and many of the short “sentences” are in fact titles and extended proper names (e.g., e-mail addresses and institution names). We excluded punctuation tokens. We also excluded tokens whose surprisal value is greater than 20 bits, as the surprisal estimates for such rare word sequences may be unreliable. Moreover, exceedingly surprising information may introduce confounding factors in human processing. We then extracted all the dependencies in which both the head and the dependent are spared from data exclusion.
3.1.5 Data Analysis
For each language, the analysis consists of three parts. The first one is on the full dataset obtained as introduced above, with all types of dependency relations included. In addition, we also took a closer look into the dependencies whose dependent is a core argument in the sentence. Therefore, we also ran analysis on two subsets of the full dataset above, which include subject relations [^5] Annotated as nsubj and csubj in UD corpora. and object relations [^6] Annotated as obj, iobj, ccomp, and xcomp in UD corpora. respectively.
For the analysis with the full dataset, for each language, we ran separate linear mixed-effects models predicting the two variants of dependency length $L$ as the dependent variable, using the lmerTest package in R (Kuznetsova et al., 2017). The critical fixed-effect predictor is the antecedent surprisal, with random intercept by dependency types. [^7] As mentioned above, we also compare the effect of antecedent surprisal with antecedent frequency in the current analysis. However, the models with random slopes for both effects rarely converge. Therefore, for better interpretability of the statistical result, we only included random intercept by dependency types. For the analyses with subject and object relations, we ran linear models with the same fixed effects. We included four control variables for all the analyses, as in (5). Sentence position aims to control the discourse-level information structure, where more information may be given as the discourse develops; Antecedent position aims to control that antecedents appearing towards to the end of a sentence naturally tend to have shorter dependency length. Sentence length aims to control for two possible confounds: first, longer sentences may tend to have longer dependency length in general; second, longer sentences may tend to have more complex syntactic structure, which may be associated with more surprising antecedents. We also included antecedent frequency in log scale retrieved from Speer (2022) in order to see whether the surprisal effect is reducible to a simple frequency effect as mentioned above. All variables are $z$-scaled.
  1. (5)
    Control variables in Study 1
    • sentence position: position of the sentence in the current document, if the corpus is organized document-by-document
    • antecedent position: position of the antecedent in the current sentence
    • sentence length: length of the sentence measured as word counts
    • antecedent frequency: log frequency of the left codependent
3.2 Result
Figure 3: Dependency length $L$ as a function of antecedent surprisal. Panel A corresponds to $L$ measured as intervening word counts. Panel B corresponds to $L$ measured as the sum of surprisal over intervening words. Surprisal is binned into 20 categories, and the mean $L$ within each category is shown with a 95% confidence interval. A linear fit to these points is presented.
Figure 4: Study 1 coefficient estimates for the effects of antecedent surprisal and antecedent frequency on dependency length $L$ across languages, with 95% confidence interval. Our hypothesis of strategic resource allocation predicts that more surprising antecedents are associated with longer $L$ (positive effect of antecedent surprisal), and that more frequent antecedents are associated with shorter $L$ (negative effect of antecedent frequency). Significance levels: *($p$$<$0.05), **($p$$<$0.01), ***($p$$<$0.001)
The result of the raw data in its original scale is presented in Figure 3, which shows dependency length $L$ as a function of antecedent surprisal. The statistical result of the regression models is presented in Figure 4 for the effects of antecedent surprisal and antecedent frequency. A high-level summary of the statistical evidence across languages is shown in Figure 5 for the critical effects. Since we used the surprisal measure from GPT-3 and mGPT, we describe an effect as robust and independent of model parameterization if it is significant in the same direction both in the analysis with GPT-3 and in the one with mGPT (i.e., both positive or both negative, highlighted in dark red and dark blue in Figure 5). [^8] Since Mandarin is not available for mGPT, a critical effect is highlighted in dark red or blue in Figure 5 even though we only have the result with GPT-3. We describe the effect as partially confirmed and less robust if it reaches significance with only one of the language models (highlighted in light red and blue in Figure 5). We describe the effect as inconclusive if it is not significant with any model, or if GPT-3 and mGPT show significantly conflicting result (i.e., significant positive in one model but significant negative in the other). [^9] The use of these terms (i.e., robust, less robust, partially confirmed, and inconclusive) is only for expository purpose, and does not imply any direct statistical test.
Figure 5: Study 1 summary of statistical result for the effects of antecedent surprisal and antecedent frequency on dependency length $L$. Significant ($p$$<$0.05) positive effects are highlighted in red; significant negative effects are highlighted in blue. Effects are considered not conclusive if insignificant with both language models, or if GPT-3 and mGPT show conflicting result where the effect is significant in opposite directions. According to our hypothesis, antecedent surprisal is expected to have a positive effect on $L$, whereas antecedent frequency is expected to show negative effect.
3.2.1 All Types of Relations
Antecedent Surprisal.
In the analysis of the full dataset with all types of dependency relations, we indeed found a significant positive effect of antecedent surprisal for the majority of the languages, whereby more surprising antecedents are associated with longer dependency length. When dependency length $L$ is measured as intervening word counts, there is a robust positive effect in six languages, namely Danish, English, German, Italian, Russian, and Spanish. When $L$ is measured as intervening surprisal, in addition to these six languages there is Mandarin that also shows a significant positive effect. However, for Japanese, Korean and Turkish, contrary to our prediction, there is a negative antecedent surprisal effect, although this unexpected negative effect is less robust for the information-theoretic $L$.
Antecedent Frequency.
The result is quite mixed for the effect of antecedent frequency. When $L$ is measured as word counts, there is a negative effect of antecedent frequency on $L$ in six languages (robust for English, Italian, Russian, and Spanish, while partially supported in Korean and Turkish). The effect, however, is unexpectedly positive in German, Japanese, and Mandarin, and is inconclusive in Danish. When $L$ is measured as surprisal, the antecedent frequency effect unexpectedly turns out to be positive for more languages, namely Danish, English, German, Italian, Japanese and Mandarin (six languages). There is still a negative effect for Spanish and Turkish, although the effect is less robust compared to when $L$ is measured as word counts.
3.2.2 Subject Relations
Antecedent Surprisal.
For subject relations, when $L$ is measured as word counts, there is a positive effect of antecedent surprisal on $L$ as predicted in five languages (English, Italian, Russian, Spanish, and Mandarin), suggesting that more surprising antecedents are associated with longer $L$ in these languages. In contrast, contrary to our prediction, the effect is negative for German, Japanese, Korean, and Turkish, while Danish shows inconclusive result. When $L$ is measured as intervening surprisal, there are seven languages that exhibit a positive antecedent surprisal effect on $L$ (English, Italian, Russian, Spanish, Japanese, Korean, and Mandarin), representing the majority of languages in our dataset. However, the effect remains negative in German, and is inconclusive for Danish and Turkish.
Antecedent Frequency.
When $L$ is measured as word counts, there is a negative effect of antecedent frequency on $L$ in six languages (Danish, English, Italian, Russian, Spanish, and Korean), suggesting that more frequent antecedents lead to shorter $L$ in these languages. The effect is unexpectedly positive in German, and is inconclusive for Japanese, Mandarin, and Turkish. When $L$ is measured as surprisal, the result is similar to the orthographic $L$, except that the negative effect in Italian and Spanish becomes less robust and the effect in Korean becomes inconclusive.
3.2.3 Object Relations
Antecedent Surprisal.
The result for object relations shows a mixed pattern. When $L$ is measured as word counts, no positive effect of antecedent surprisal on $L$ is observed in any language. Instead, there is a negative effect in German, Russian, Spanish, Korean, and Mandarin, and the result is inconclusive for the rest of the languages. However, the pattern is reversed for the information-theoretic $L$, with a positive antecedent surprisal effect being observed in seven languages (Danish, English, Italian, Russian, Spanish, Japanese, Turkish). Only in Korean did we observe a negative effect, and the result is inconclusive in German and Mandarin.
Antecedent Frequency.
When $L$ is measured as word counts, we only found a robust antecedent frequency effect on $L$ in four languages, two negative (Italian and Spanish) and two positive (English and German). The result for the rest of the languages is inconclusive. When $L$ is measured as intervening surprisal, there are more languages that show an unexpectedly positive effect (Danish, English, German, Italian, Korean, and Mandarin). Only in Spanish did we observe a negative antecedent frequency effect. The result for Russian, Japanese and Turkish remains inconclusive.
3.3 Discussion
In this cross-linguistic corpus study, we indeed found evidence for a positive effect of antecedent surprisal on dependency length $L$, especially when $L$ is measured as intervening surprisal (bottom left panel in Figure 5), which, as mentioned above, should better capture the variability of information load across word types. With this information-theoretic dependency length measure, the majority of languages (seven out of ten) exhibit a positive effect of antecedent surprisal on $L$, and this effect still holds when we zoom into the subset that only includes subject or object relations. Overall, as predicted, this pattern indicates that more surprising antecedents are associated with longer dependency length, suggesting that the pressure to minimize dependency length is relaxed when the antecedent is of higher surprisal. Consistent with our hypothesis of strategic resource allocation, the result supports that novel and unexpected linguistic units can tolerate longer dependency length before its retrieval site, possibly because unexpected information is prioritized for working memory resources during encoding, and is more resistant to memory decay and interference.
However, there are two caveats worth noting. First, there is considerable cross-linguistic variability in our result, and the antecedent surprisal effect mostly exists within the Indo-European and head-initial languages in our analysis. Second, although the analysis on the full dataset with all types of dependencies reveals a general trend for a positive antecedent surprisal effect, the result is much more consistent within subject relations across different dependency length measures than within object relations. In object relations, the expected effect is reversed with the orthographic dependency length.
It is also worth noting that the positive antecedent surprisal effect on $L$ cannot be reduced to a pure frequency effect. That is, there is still a significant effect of antecedent surprisal even though antecedent frequency has been included in the regression models as a control variable. Moreover, compared to antecedent surprisal, the effect of antecedent frequency on $L$ is much less consistent, and better aligns with our prediction with the orthographic $L$. That is, more frequent antecedents are associated with shorter $L$ mostly when $L$ is measured as word counts instead of intervening surprisal. Taken together, the antecedent surprisal effect we found on dependency length is irreducible to a pure frequency effect, and the antecedent frequency effect by itself is less consistent.
In the end, to what extent does written corpus text approximate language production? Compared to spoken language, written language typically allows “speakers” more time to think, reducing much of the cognitive load involved in production, and the communicative goal is more geared towards listeners’ need. That being said, speaker-oriented cognitive constraints, such as memory capacity, may play a less prominent role, and the need for strategic memory allocation may be diminished in written language production. Therefore, the effect observed in the current analysis using written text can be viewed as an upper bound, and we expect the effect of strategic memory allocation to be stronger when using spoken language corpora.
4 Study 2a: Comprehension Side (Self-Paced Reading)
In this second study, we investigate whether the effect of strategic resource allocation also holds from the comprehension side. In particular, we examine to what extent the dependency locality effect observed in previous studies (Gibson, 1998, 2000; Grodner & Gibson, 2005; Bartek et al., 2011) can be modulated by the surprisal of the antecedent. First, we expect a baseline dependency locality effect, where the processing difficulty at the retrieval site, manifested as reading time, is expected to increase as the dependency length gets longer. Second, according to our hypothesis, more surprising antecedents are more capable of tolerating stronger memory interference. That is, the longer dependency length does not create too much additional processing difficulty for the retrieval of surprising antecedents, resulting in a reduced locality effect. Consider the example below in (6):
  1. (6)
    1. a.
      The evil monster{ $\text{The evil {monster}}\begin{Bmatrix}\text{ in the tower }\\ \text{ who stayed in the tower }\\ \text{ who stayed in the tower near the castle }\end{Bmatrix}\text{\text@underline{approached}...}$
    2. b.
      The cute monster{ $\text{The cute {monster}}\begin{Bmatrix}\text{ in the tower }\\ \text{ who stayed in the tower }\\ \text{ who stayed in the tower near the castle }\end{Bmatrix}\text{\text@underline{approached}...}$
As explained in Study 1, “the cute monster” in (6b) is more prioritized with encoding resources for its lower predictability, making it capable of tolerating longer dependency length. Therefore, when there is longer dependency distance, the processing difficulty at the retrieval site (i.e., the main verb “approached” in this example) should increase at slower rate in (6b) than in (6a), since the additional intervening material induces lower level of interference for the more surprising antecedent in (6b). As a result, on top of the baseline effect on retrieval difficulty, we expect a negative interaction between dependency length $L$ and antecedent surprisal at the retrieval site.
4.1 Method
4.1.1 Data
The data we used in Study 2a is taken from the Natural Stories Corpus (NSC) (Futrell et al., 2021). The text of the corpus is in English, and contains 10,245 lexical words in 485 sentences, taken from 10 stories with around 1000 words each. The corpus includes dependency annotation. The reading time (RT) data is collected from 181 native English speakers, using the self-paced reading task (SPR). The original corpus already excluded participants with low comprehension accuracy, as well as the reading times either faster than 100ms or slower than 3000ms. Therefore, we did not perform additional exclusion of reading time data in the current study. We generated the surprisal estimates for each word from mGPT. [^10] GPT-3 is no longer accessible from OpenAI since January 4th, 2024. Therefore, we only used the surprisal estimates from mGPT for the analyses in Study 2a and 2b. The NSC corpus comes with dependency annotation in UD style.
4.1.2 Data Transformation, Exclusion, and Analysis
As in Study 1, we analyzed three RT datasets as well here in the current study, namely the full dataset with all types of dependencies, a subset with subject relations only, and a subset with object relations only. The sample size is summarized in Table 2. We ran linear mixed-effect models on the log-transformed RTs at the right codependent of each dependency, which is considered the retrieval site for the antecedent. The critical effect is the interaction between dependency length $L$ and antecedent surprisal, with maximal converging random intercept and random slopes by participant. For the analysis of the full dataset, we also included maximal converging random effects by dependency type.
Study 2b
Study 2a First-pass Total RT
All types 601,122 256,567 313,167
Subject relations 57,023 21,709 26,885
Object relations 48,091 20,361 24,990
Table 2: RT data sample size in Study 2a and 2b
The four control variables applied in Study 1 (namely, sentence position, antecedent position, sentence length, and antecedent frequency) are also included here in Study 2a. Besides these, we included several additional control variables that are often considered highly relevant for reading time measures. First, we included word length, word surprisal and word frequency of the right codependent itself. Second, to control the spillover effect often found in reading studies, we included word surprisal and word frequency of the two previous words before the right codependent. As in Study 1, frequency measures are in log scale, generated from Speer (2022). All variables are $z$-scaled. The transformation and exclusion of the dependency data follow the same procedure as in Study 1.
4.2 Results
Figure 6: Study 2a reading time (RT) result. Panel A: raw RTs at the right codependent as a function of dependency length $L$ modulated by antecedent surprisal, which is binned into tertiles for visualization. Panel B: result of regression models with log-transformed RTs; coefficient estimates with 95% confidence interval for the effects of $L$, antecedent surprisal, and the interaction between the two. Significance levels: *(p$<$0.05), **(p$<$0.01), ***(p$<$0.001)
Figure 6A shows the interaction effect between dependency length $L$ and antecedent surprisal on raw reading times. The result of statistical models for the critical effects is summarized in Figure 6B.
All types of relations.
In the analysis of the full dataset with all types of dependency relations, we first found a baseline locality effect, where dependency length $L$ has a positive main effect on RTs at the retrieval site (i.e., the right codependent), suggesting that longer distance between codependents makes it more difficult to retrieve the antecedent at the right codependent. This baseline locality effect holds for both measures of $L$. However, there is no evidence for an antecedent surprisal main effect. In the end, importantly, we found a negative $L$$\times$ antecedent surprisal two-way interaction. Consistent with our prediction, this negative interaction suggests that the locality effect on the RT of right codependents is reduced when the antecedent is more surprising. This two-way interaction, however, only holds when $L$ is measured as surprisal, and no interaction is found when $L$ is measured as word counts.
Subject relations.
First, for both measures of dependency length $L$, a baseline locality main effect of $L$ was found for subject relations, where longer $L$ leads to longer RT at the right codependent. Second, as in the analysis of the full dataset above, there is no antecedent surprisal main effect with either measure of $L$. In the end, we found a critical negative $L$$\times$ antecedent surprisal two-way interaction with both $L$ measures. As predicted, this negative interaction is indicative of a reduced locality effect for more surprising antecedents.
Object relations.
The result of object relations has a relatively complex pattern. First, unlike the previous two analyses, there is an unexpected negative, not positive, main effect of $L$ on the RT at retrieval site. This negative main effect reaches significance when $L$ is measured as surprisal, and still numerically holds when $L$ is measured as word counts. Instead of a baseline locality effect, this negative $L$ effect suggests an anti-locality effect, where more intervening material between the two codependents leads to faster RT at the retrieval site. Moreover, as seen in Figure 6A (right column), this anti-locality effect is mostly driven by antecedents of low surprisal. Second, there is a negative antecedent surprisal main effect for both $L$ measures, whereby more surprising antecedents induce faster RT at the retrieval site. In the end, unlike the negative $L$$\times$ antecedent surprisal interaction observed in the previous two analyses, object relations exhibit a positive $L$$\times$ antecedent surprisal interaction, which is significant for both $L$ measures. However, since the main effect of $L$ is negative in the first place, instead of an enhanced locality effect, this positive interaction actually suggests a reduced anti-locality effect for more surprising antecedents. It is not yet entirely clear to us why there is an anti-locality effect in the first place exclusively for object relations, but the result pattern in object relations seems to point to a potential trade-off between the direction of the $L$ main effect and the direction of the $L$$\times$ antecedent surprisal interaction, which we will discuss below in Section 4.3.
4.3 Discussion
To sum up, in Study 2a, through the analysis of the RT data in the Natural Stories Corpus, we first replicated the baseline locality effect in the analysis of all types of dependency relations and in the one of subject relations only, suggesting that the non-local retrieval of the antecedent at the right codependent becomes more difficult when there is more intervening material. However, we also found an anti-locality effect in object relations, which suggests that more intervening material actually facilitates the establishment of a non-local object relation. Second, only in object relations did we observe a negative main effect of antecedent surprisal on the RT at the right codependent, suggesting that more surprising antecedents are easier to retrieve in object relations. This negative effect of antecedent surprisal is consistent with the findings in Hofmeister (2011), which observes that semantically more complex noun phrases are easier to be retrieved. However, more future work is needed to investigate why this effect only emerges in object relations. In the end, the most important finding of this Study 2a is the interaction between dependency length and antecedent surprisal. Specifically, in the analysis of the full dataset and the one of subject relations, we found a reduced locality effect for more surprising antecedents. This reduced locality effect is consistent with the prediction of strategic resource allocation, in that more surprising antecedents are prioritized for working memory resources and are encoded with more robust representation against memory interference and decay.
It is also worth noting that there seems to be a potential trade-off between the direction of the $L$ main effect and the direction of the $L$$\times$ antecedent surprisal interaction. On the one hand, when the main effect of $L$ is positive, as in the first two analyses, there is a negative $L$$\times$ antecedent surprisal interaction, suggesting a reduced locality effect for more surprising antecedents. On the other hand, when the main effect of $L$ is negative to start with, as in the object relations, the interaction becomes positive, indicating that the anti-locality effect is reduced for more surprising antecedents. The reduced locality effect is straightforward, as predicted by our hypothesis. But why is there a reduced anti-locality effect? In fact, the reduced anti-locality effect for more surprising antecedents can be consistent with our strategic resource allocation as well. According to experience-based processing theories, the anti-locality effect can be viewed as a facilitation effect on the prediction of the right codependent. That is, more intervening material may provide more information about the identity of the word at the right codependent, helping the comprehender to make better predictions (Levy, 2008a), thus canceling out the burden created by memory interference. However, for more surprising antecedents, if their representation is more enhanced due to strategic resource allocation, it is possible that comprehenders can already rely on the their memory of the antecedent to predict the right codependent. As a result, the intervening material may no longer provide too much additional help to make predictions. This reduced facilitation from the intervening material for more surprising antecedent, therefore, may manifest itself as a reduced anti-locality effect.
5 Study 2b: Comprehension Side (Eye-Tracking)
In Study 2b, we examine strategic resource allocation in dependency locality using an eye-tracking corpus. As in Study 2a, we expect to see a baseline locality effect, as well as an interaction between locality and antecedent surprisal, in the sense that a reduced locality effect is associated with more surprising antecedents.
5.1 Method
5.1.1 Data
The data we used in Study 2b is taken from the English part of Dundee corpus (Kennedy & Pynte, 2005). The corpus consists of 20 texts, with 56,212 tokens in total (around 2800 words for each text). The eye-tracking data is collected from 10 English native speakers, with each text being split into 40 fine-line screens. We analyzed two eye-tracking measures: first-pass reading time, defined as the sum of all the fixations on a region after first entering in the region and before first leaving it either to the left or to the right; and total reading time, defined as the sum of all the fixations on a region throughout a trial. Like NSC corpus, the Dundee corpus also comes with UD-style dependency annotation.
5.1.2 Data Transformation, Exclusion, and Analysis
The transformation and exclusion of dependency data follow the same procedure as in Study 1 and Study 2a. The reading time responses are excluded if shorter than 100ms or longer than 3000ms. The sample size after data exclusion is summarized in Table 2. We ran linear mixed-effects models on first-pass durations and total reading times. Both reading time measures are log-transformed. As in Study 2a, the critical effect is the interaction between dependency length $L$ and antecedent surprisal, with the same random structure and control variables as in Study 2a.
5.2 Results
Figure 7: Study 2b reading time (RT) result for first-pass duration and total RT. Panel A: raw RTs at the right codependent as a function of dependency length $L$ modulated by antecedent surprisal, which is binned into tertiles for visualization. Panel B: result of regression models with log-transformed RTs; coefficient estimates with 95% confidence interval for the effects of $L$, antecedent surprisal, and the interaction between the two. Significance levels: *(p$<$0.05), **(p$<$0.01), ***(p$<$0.001)
Figure 7A shows the interaction effect between dependency length $L$ and antecedent surprisal on raw reading times, including both the first-pass duration and the total RT. The result of statistical models for the critical effects is summarized in Figure 7B.
All types of relations.
As shown in Figure 7B (left column), no evidence was found for the main effect of dependency length $L$ with any measure of RT and $L$, suggesting the lack of the baseline locality effect. No significant main effect of antecedent surprisal, either. For the critical $L$$\times$ antecedent surprisal interaction, although the effect is numerically negative on first-pass RT with both measures of $L$, it is only marginally significant. No evidence for the interaction effect was found on total RT.
Subject relations.
Still, as shown in Figure 7B (middle column), we did not find any evidence for either an $L$ or an antecedent surprisal main effect with any measure of RT and $L$. However, there is indeed a negative $L$$\times$ antecedent surprisal two-way interaction, suggesting that the locality effect at the retrieval site, although not statistically significant on average, is reduced for more surprising antecedents. This interaction effect reliably holds for first-pass RT with both $L$ measures, as well as for total RT with $L$ measured as word counts. It is, however, only marginally significant for total RT with $L$ as intervening surprisal.
Object relations.
As shown in Figure 7B (right column), we found in object relations a baseline locality effect manifested as a positive $L$ main effect on RTs at the retrieval site, which holds for first-pass RT with $L$ as word counts and for total RT with both $L$ measures. There is no evidence for an antecedent surprisal main effect. In terms of the critical $L$$\times$ antecedent surprisal interaction, we only found a marginally significant negative interaction for total RT with $L$ measured as surprisal, and the effect is non-significant elsewhere.
5.3 Discussion
The result of the two main effects in the current Study 2b shows a very different pattern compared to Study 2a. First, unlike Study 2a, in the current Study 2b only in object relations did we find a baseline locality effect, whereas the locality effect is not observed in the analysis of the full dataset or in the one of subject relations. That is, longer dependency length does not lead to higher processing difficulty at the retrieval site for subject relations. This lack of the baseline locality effect aligns with the observation in Demberg and Keller (2008), where the locality effect in Dundee corpus is overall small and unreliable for verbs, which in our case is the retrieval site of subject relations. Second, there is no main effect of antecedent surprisal across the board in this Study 2b, casting doubt on how reliable the negative antecedent surprisal main effect is in Study 2a.
Although the baseline locality effect is relatively unreliable, we still observed evidence for a negative $L$$\times$ antecedent surprisal interaction effect in this Study 2b, especially in subject relations. As shown in Figure 7A (middle column), compared to antecedents with low-to-mid surprisal levels, those with high surprisal exhibit weaker locality effect. Similar to the interaction effect observed in Study 2a, the current result shows that more surprising antecedents is less susceptible to the effect of memory interference induced by intervening material, possibly due to their enhanced representation. Supplementing the self-paced reading data in Study 2a, the result of Study 2b thus lends support to our hypothesis of strategic resource allocation with data from eye-tracking paradigm.
6 General Discussion
In three corpus studies, we examined strategic resource allocation through the lens of dependency locality both in production and in comprehension. Study 1 explored this hypothesis in production by analyzing UD corpora from 10 languages. Our result reveals that more surprising antecedents can tolerate more intervening material before they need to be retrieved at the other side of the dependency, resulting in a positive correlation between antecedent surprisal and dependency length. However, it is worth noting that this reduced locality effect mostly exists within Indo-European and head-initial languages, and is more consistent for dependencies of subject relations than for object relations.
In Study 2, we shifted gears and focused on comprehension, analyzing two English reading-time corpora: one based on self-paced reading paradigm (Study 2a) and the other based on eye-tracking paradigm (Study 2b). The SPR data reveals a baseline locality effect, where longer dependency lengths lead to increased reading times at the retrieval site. Importantly, we found a $L$$\times$ antecedent surprisal interaction, where the baseline locality effect is reduced for more surprising antecedents. Moreover, we again observed a subject-object asymmetry, such that the critical effect consistently holds only in subject relations. The eye-tracking data shows a more nuanced pattern: although the baseline locality effect was observed only in object relations, we indeed found an $L$$\times$ antecedent surprisal interaction in subject relations.
Overall, despite the caveats mentioned above, our result shows converging evidence that a reduced locality effect emerges for more surprising antecedents in the processing of non-local dependencies, suggesting that more surprising antecedents are less susceptible to the interference from intervening material. This finding aligns with the notion of strategic resource allocation that we proposed, which holds that unexpected information is prioritized for memory resources and is encoded with enhanced memory representation.
6.1 The role of context in working memory efficiency
Earlier in Introduction, we have argued that the efficiency of working memory allocation depends on how likely a linguistic unit can be reconstructed based on the statistical structure of linguistic input. But an open question is: what kinds of statistics are being used? More specifically, to what extent is working memory efficiency guided by context-sensitive statistics?
The role of contextual information has been under debate across multiple domains of linguistic efficiency, especially in the context of signal reduction (Jaeger & Buz, 2017). One of the earliest evidence supporting the communicative efficiency pressure for linguistic structure is the well-known Zipf’s law, which observes that more frequent words tend to have short forms (Zipf, 1949). Similar reduction effect of linguistic forms as a function of usage frequency has also been found in the historical change of linguistic representations (e.g., Bybee et al., 1994; Bybee, 2006; Pierrehumbert, 2008; Cohen Priva, 2015). While most of these theories focus on frequency, which can be considered unigram probabilities independently generated from a stationary distribution, recent studies have begun investigating the role of context-specific probabilities in linguistic efficiency. The findings in this area are mixed. For example, building on Zipf’s observation, Piantadosi et al. (2011) find that a significant amount of word-length variability is explained by contextual predictability in addition to the frequency effect. In contrast, Pimentel et al. (2023) argue that word length is better predicted by frequency. Beyond the structure of lexicons, in online processing, contextual predictability also shapes the reduction of referring expressions (e.g., Mahowald et al., 2013; Tily & Piantadosi, 2009; Xu & Xiang, 2021) as well as syntactic structure (e.g., Jaeger, 2010; Jaeger & Levy, 2006).
In the current study, that the observed antecedent surprisal effect cannot be reduced to a pure frequency effect suggests that the statistics relevant to working memory efficiency go beyond unigram frequencies, and the strategic allocation of working memory resources is based on more fine-grained context-specific probabilities. As noted by Jaeger & Buz (2017), frequency can be understood as an averaged effect of contextual predictability. This is probably one of the reasons why linguistic theories focusing on the representational aspects of language often emphasize frequency, as it reflects the abstract, global properties of a language accumulated through long-term experience. That being said, the optimization of working memory efficiency not only relies on those ready-to-retrieve statistics already stored in long-term memory, but also incorporates statistics computed online, dynamically drawing upon rich contextual information in a rapid and adaptive manner (for a different view, see Opedal et al., 2024). It is worth noting that by using surprisals from autoregressive language models the current study only tests the left context of a linguistic unit. A natural direction for future work is to explore the role of bidirectional context in working memory efficiency.
6.2 Relationship with lossy-context surprisal
The mechanism of strategic resource allocation we proposed bears similar theoretical and empirical implications with lossy-context surprisal theory and its variants (Futrell, Gibson, and Levy, 2020; Hahn et al., 2022). Focusing on the prediction mechanism, lossy-context surprisal holds that next-word predictions are based on lossy and faulty memory representations, rather than the veridical representation of the past linguistic input. The theory explicitly includes a memory distortion process, where certain elements in an utterance are erased to form a lossy representation, subject to certain probability distributions.
In many aspects, our proposal and lossy-context surprisal form two sides of the same coin. For us, we seek to understand and explain the working memory mechanism in sentence processing, and the mechanism we proposed includes a component of predictive processing. For lossy-context surprisal, the theory aims to understand the prediction mechanism in sentence processing, which includes a memory component. However, we both speak for an interaction between memory and prediction, as both of them should jointly support human linguistic behaviors as a cognitive task. That means, our proposal and lossy-context surprisal are not mutually exclusive, theoretically or empirically, and we simply focus on different perspectives of a similar cognitive task.
In fact, our theory of strategic resource allocation and lossy-context surprisal are on some level mutually translatable, so we do not see them necessarily as conflicting theories. On the one hand, in the language of lossy-context surprisal, the memory distortion process where certain linguistic units are erased is potentially where our strategic resource allocation could fit in, such that more surprising units given the context are more likely to be erased when predicting future units. On the other hand, the memory component in lossy-context surprisal also fits in the resource-rational explanation of working memory efficiency. As outlined above in Introduction, a fundamental question at the core of our resource-rational analysis of working memory mechanism is: if memory capacity is limited, how to minimize the distortion cost by strategically prioritizing more important linguistic units? However, what counts as important? Or, in other words, what is the objective function based on which the distortion cost is defined? In the original lossy-context surprisal theory (Futrell, Gibson, and Levy, 2020), this objective function is not explicitly specified. In Hahn et al. (2022), the model takes one step further, and this objective function is to minimize the downstream next-word prediction task. In our proposal, the distortion cost is defined by how likely a unit can be reconstructed later given the context. These two different objective functions of distortion cost are not mutually exclusive, and the accuracy of next-word prediction may be compromised if a lost unit in memory is not reconstructable. Therefore, although our proposal of strategic resource allocation has a different theoretical focus from the model of Hahn et al. (2022), our empirical predictions actually share some overlap, and we both predict on some level that the representation of more surprising units (or, less frequeny units) should be more robust (or, less likely to be deleted in the lossy-context surprisal model).
6.3 Hierarchical encoding and compression
Strategic resource allocation, arising as an efficiency principle from the functional pressures of working memory, is situated more at the computational level in Marr (1982)’s three-level representation. A natural question to ask then is: what is the potential mechanism to implement this efficiency principle at the algorithmic level?
One possible mechanism is hierarchical compression, which postulates that information can be stored in memory with a multi-level hierarchy of abstraction (Brady et al., 2009; Bates and Jacobs, 2020; Brady et al., 2024; Craik and Lockhart, 1972; Christiansen and Chater, 2016). In visual working memory, higher levels are more compressed, having a more categorical nature; lower levels, in contrast, are encoded with more quantitative perceptual detail. A similar hierarchy also exists in sentence processing. Sequential linguistic input can be continuously encoded and recoded into compressed forms, which in turn are further compressed into more abstract forms when new input comes in (Christiansen and Chater, 2016). This incremental compression procedure gives rise to a multi-level hierarchical structure of memory representation, such that higher levels of abstraction are encoded as a gist of message without specifying elaborated syntactic and semantic features (Bradshaw and Anderson, 1982). Intuitively, more memory resources should yield more detailed encoding. Indeed, this has been recently demonstrated by some memory encoding models grounded in the rate-distortion theory, where a quantitative-categorical spectrum naturally arises simply by manipulating the memory capacity during encoding (Bates and Jacobs, 2020; Jakob & Gershman, 2023).
Predictable information, if less prioritized, should be encoded in a more compressed and abstract fashion. In visual working memory, this is indeed evidenced by the fact that more items in a visual memory task can be stored when their perceptual features are statistically correlated (Brady et al., 2009; Bates et al., 2019). Moreover, information with stronger prior knowledge is also more susceptible to the categorical bias in perception, where the perceived input is biased towards the categorical mean (Bates and Jacobs, 2020). In the domain of language, more frequent linguistic sequence is more likely to be holistically stored in memory (e.g., Bybee, 2006; Goldberg, 2003; Hawkins, 2004; Traugott & Trousdale, 2013). Importantly, hierarchical compression based on statistical regularities provides a mechanism that strings together effects across different linguistic representational levels, forming a spectrum of compression. At one end, there is the locality effect where words that are more mutually predictable tend to stay closer to each other in linear order (e.g., Futrell et al., 2019; Futrell, 2019). At the other end, the same pressure of compression governs the fusion of morphemes (e.g., Hahn et al., 2021; Rathi et al., 2021), and makes mutually predictable units more likely to go through processes such as affixation and phonological reduction (e.g., Bybee et al., 1994; Bybee, 2006; Gahl and Baayen, 2024).
6.4 Parallelism between production and comprehension
The effect of strategic resource allocation holds both for production and for comprehension in the current study, pointing to the parallelism between these two modalities. For comprehension, a reasonable question is: to what extent the observed effect of strategic resource allocation is experience-based (MacDonald & Christiansen, 2002), given that the same effect is also seen in production data? In other words, it is possible that comprehenders prioritize unexpected linguistic units and encode them with enhanced representation because unexpected units are more likely to associate with stronger memory interference in the production data they receive.
A similar question can also be asked for production: to what extent the effect observed in production is the result of audience design (H. H. Clark and Murphy, 1982; Ferreira, 2019; Lockridge & Brennan, 2002), given that comprehenders can encode unexpected units with prioritized memory resources. One listner-oriented production theory compatible with our finding is the Uniform Information Density (UID) theory (Jaeger & Levy, 2006; Meister et al., 2021; T. H. Clark et al., 2023). According to UID, surprising antecedents may be followed by longer dependency length so that there is a smoother transition to the other side of the dependency, which is relatively predictable since it shares high mutual information with the antecedent (Futrell et al., 2019).
Here we do not attempt to adjudicate the possibilities mentioned above, nor do we view these possibilities in conflict with our proposal. In our opinion, strategic resource allocation provides a potential explanation for the mechanistic underpinnings of these higher-level processes.
Despite the parallelism, comprehension and production still differ in some critical aspects, exerting modality-specific constraints on strategic resource allocation due to their idiosyncratic processing nature. For example, production is, in general,l, more cognitively demanding, in need of higher memory capacity, executive control, and action planning than comprehension (MacDonald, 2013; Nozari & Novick, 2017; Hickok, 2012; Koranda et al., 2020). This additional cognitive demand may exert more pressure to efficiently use the limited memory resources in production than in comprehension, possibly resulting in a stronger effect of strategic resource allocation. Future work is needed to investigate this possibility.
6.5 Implications for linguistic typology
Dependency length minimization as a functional universal has been argued to shape the syntactic structure of human language. As reviewed in Introduction, this universal arises from the pressure to efficiently use the limited working memory resources. In the current study, we go beyond the general constraint of limited memory capacity, and further argue that memory resources should be strategically allocated to prioritize novel and unexpected information, a memory efficiency principle that naturally arises from two assumptions about working memory. The results indicate that this strategic resource allocation indeed serves as a functional constraint to shape syntactic structures: although there is a general pressure to minimize dependency length, this pressure can actually be relaxed when the antecedent of a syntactic dependency is of higher surprisal. Our finding further substantiates the functionalist view as a promising approach to provide explanatory accounts for linguistic universals (Gibson et al., 2019). It also highlights the importance of having increasingly sophisticated characterization of functional constraints, in order to see how far this functional perspective of language structure can go.
How does strategic resource allocation as a universal efficiency principle interact with other grammatical constraints and language-specific phrase structures? One consistent pattern we observed is the asymmetry between subject and object relations. Specifically, the effect of strategic resource allocation is generally smaller and less reliable for object relations than for subject relations. One possible explanation is that object relations are subject to stronger grammatical constraints, with greater pressure to position the head and its dependent closer to each other. In support of this possibility, Keenan & Comrie (1977) identify an Accessibility Hierarchy as a linguistic universal, where noun phrases in the subject position are more readily relativized into relative clauses than those in the object position. Such constraints may bind object noun phrases and verbs more tightly, reducing the influence of strategic resource allocation. Moreover, in many languages, grammatical agreement is common in subject-verb dependencies but absent in object-verb dependencies. Therefore, establishing subject relations may require more active grammatical computation, resulting in increased memory demand in processing subject relations and a stronger need for more strategic and efficient use of working memory resources.
Another notable pattern in Study 1 is that, compared to head-initial languages, most of the head-final languages in our analysis do not exhibit a reliable antecedent surprisal effect on dependency length. We speculate that this may be related to the tendency for argument dropping in head-final languages with SOV word order. From the perspective of dependency locality, SOV word order is associated with longer dependency lengths compared to SVO, which should theoretically impose higher memory costs and make it a less efficient structure. Despite this seemingly inefficiency, typologically, SOV is a word order commonly attested (Hammarström, 2016). To address this paradox, some studies find that arguments in SOV languages are often dropped, thus reducing the overall dependency length in actual language use (Levshina, 2025; Ueno & Polinsky, 2009). It is possible that by allowing speakers to drop arguments, the efficiency of SOV structure may already be significantly improved, obviating the need for strategic resource allocation as another efficiency strategy.
It is also worth mentioning a few other relevant language-specific factors. For Mandarin Chinese, although it predominantly follows SVO word order, the relative clause goes before the nouns, increasing the flexibility of object relations in terms of their dependency length. For Korean and Japanese, the subject noun phrase may be delayed when it is surprising and unexpected, shifting the word order from SOV to OSV, and therefore decreasing the dependency length of subject relations for surprising antecedents.
6.6 Surprisal effect as efficient coding
One way to interpret our finding is that the enhanced robustness of memory representation arises as a by-product of the effort involved in processing surprising information. However, this raises an even more fundamental question: why does the surprisal effect occur in the first place? In other words, why is there a widely observed positive relationship between surprisal and processing effort?
Here is one way to think about the basic surprisal effect from an information-theoretic perspective. More surprising linguistic units correspond to longer code length. For example, consider a low-surprisal word encoded by a sequence of 3 bits ‘110’, compared to a high-surprisal word is encoded by 11 bits ‘11100001101’. A simple mechanical interpretation of the basic surprisal effect, therefore, is that encoding longer sequence of code in memory requires more time and effort. This leads to the widely observed linear relationship between surprisal and behavioral measures, such as reading time (e.g., Smith & Levy, 2013; Hoover et al., 2023; Shain et al., 2024; Wilcox et al., 2023; Xu et al., 2023).
However, choosing a code length based on predictability is not the only possible strategy to encode a linguistic unit into memory. An alternative approach is to assign each unit a code of equal length regardless of statistical regularities, a scheme analogous to the uniform distribution of resource allocation discussed in the Introduction. For example, the ASCII (American Standard Code for Information Interchange) system uses such a coding strategy, where every character is represented as a 7-bit sequence. Under this hypothetical scenario, processing effort would be insensitive to statistical regularities and uniformly distributed across all linguistic units.
As demonstrated in previous work on information theory, a coding scheme that treats every input as equally likely is inherently inefficient. When the statistical structure of the input is known, a coding scheme that assigns shorter code lengths to more predictable inputs can reduce the average code length, thus increasing efficiency (Shannon, 1948). The observed linear relationship between surprisal and processing effort suggests that the brain may adopt such an efficient coding strategy. Specifically, the brain appears to assign code lengths to input units in proportion to their likelihood of occurrence based on long-term statistical regularities. This strategy is supported by the evidence that the brain is very good at inferring and approximating the statistical structure of the external environment (Saffran et al., 1996), and there is good reason to believe that the brain uses these inferred statistical structure to encode information more efficiently.
The connection between the surprisal effect and efficient coding suggests that the basic surprisal effect by itself may be the result of evolutionary adaptation to efficiency pressures. Under these pressures, the cognitive system may have evolved to adopt a memory encoding strategy that optimizes an objective function within the constraints of limited resources. The exact nature of this objective function remains an open question: it could be the minimization of distortion cost, or it could involve something else. The key point, however, is that the relationship between surprisal and strategic resource allocation can be understood at different timescales. On the one hand, at the level of processing individual sentences, strategic resource allocation may arise as a by-product of the effort required to process surprising information. On the other hand, over a longer timescale, the tendency to invest more effort in encoding less predictable information may reflect an evolved strategy that is adapted to an efficiency problem.
6.7 Limitations
One major limitation of our analysis is that it heavily depends on the quality of the surprisal measures generated from LLMs. In the current study, we conducted our analysis using the surprisal from two language models, namely GPT-3 and mGPT. As presented above, the two models do yield consistent pattern, alleviating the concern that our main result may be the artifact of any model-specific behavior. However, it still does not entirely rule out the issue with the accuracy of LLM surprisals, especially for low-resource languages that are under-represented in the training data of the models we used. This limitation may compromise the extensibility of our analysis to under-studied languages, which are of particular interest from a typological perspective, and is particularly relevant for the unreliable effect we observed for some non-Indo-European languages.
In the end, even though contemporary LLMs can provide state-of-the-art probabilistic measures for linguistic data, it remains questionable to what extent it reflects the predictive processing in humans. Despite the correlation between the model-generated surprisals and the behavioral or neural responses in humans, many studies actually find that there still remain some critical patterns in human empirical data that cannot be fully accounted for solely by surprisals (Huang et al., 2024; van Schijndel & Linzen, 2021). Moreover, compared to humans, modern LLMs are far less constrained in terms of their memory capacity. This makes LLMs less likely to resemble the memory architecture in humans, or to capture the memory processes stemmed from the efficiency pressure exerted by the limited memory capacity (Oh & Schuler, 2023; Timkey & Linzen, 2023). However, this is not necessarily a limitation for the current study, since there are cases where a model with superhuman memory can provide probabilistic measures that more accurately reflect the statistical properties in the linguistic data without being confounded by the memory interference in the language model itself.
7 Conclusion
The current study investigates strategic resource allocation as an efficiency principle for memory encoding in sentence processing, which holds that working memory resources are strategically and dynamically allocated to prioritize novel and unexpected information. Empirically, this principle predicts that the memory representation of more surprising linguistic units is more robust against interference and decay. We examined this prediction through naturalistic corpus data in the context of dependency locality from both the comprehension and the production side. In production, through the analysis of UD corpora in 10 languages, we indeed found that more surprising antecedents can tolerate longer dependency length, but the effect mostly exists within Indo-European and head-initial languages. This cross-linguistic variability therefore calls for a closer look into the interaction between strategic resource allocation as a universal efficiency principle and language-specific phrase structure in future work. In comprehension, through two English reading time corpora, we observed a similar reduced locality effect on retrieval difficulty for more surprising antecedents. Moreover, in all analyses, we found that the effect is more reliable for dependencies of subject relations than object relations. Taken together, there is converging evidence from naturalistic corpus data supporting that unpredictable antecedents are encoded with enhanced representation to be more resistant against memory decay and interference, a pattern that is predicted by strategic resource allocation. For linguistic typology, our findings contribute to the functionalist approach, highlighting the need to examine more sophisticated functional constraints in terms of their explanatory power for the structure of human language.